How many moles of hydrogen are required to react with 2.5 moles of oxygen to produce hydrogen oxide? Show work.




1
Expert's answer
2022-02-09T14:35:01-0500

According to reaction below, when one mol of oxygen reacts with two mol of hydrogen, two mol of water is produced. Thus, when 2.5 mol of oxygen reacted with hydrogen, 2 * 2.5 mol = 5 mol of water is produced. The molar mass of water is 18 g/mol. One mol of water weighs 18 g. 5 mol of water weights 5 * 18 = 90 g.

O2 + 2 H2 = 2 H2O